OpenStudy (anonymous):

The rock in a lead ore deposit contains 90% \% PbS {\rm PbS} by massHow many kilograms of the rock must be processed to obtain 4.5kg {\rm kg} of Pb {\rm Pb}?

OpenStudy (cuanchi):

\[4.5kg.Pb \times\frac{MM.PbS}{MM.Pb}\times\frac{100kg.ore}{90kg.PbS}\] MM= molecular mass (calculate from the periodic table) can be in Kg or g but both in the same units